### v1

# revtex4-1Repair the float

Abstract
We report the measurement of the electric dipole moment of aluminum monochloride (AlCl) using a cryogenic buffer-gas beam source. Our measurements provide values for the dipole moments of the two lowest vibrational states of the $X^{1}\Sigma^{+}$ and the $A^{1}\Pi$ electronic states. We also show that spin-orbit coupling with an extended number of spin states is essential in the ab initio calculation to correctly describe both the dipole moment and the $T_{e}$ energy of AlCl. We further lay out the implications of these results for astrophysical models of stellar and planetary evolution that have used a substitute value for the dipole moment of AlCl until now.
The ability to control molecules and the interactions between them are expected to enable a vast number of applications McCarron et al. (2018); Fitch and Tarbutt (2021), including novel quantum computing and simulation platforms DeMille (2002); Yelin et al. (2006); Yu et al. (2019); Carr et al. (2009); Micheli et al. (2006); Bao et al. (2022); Holland et al. (2023), controlled quantum chemistry Krems (2008); Ospelkaus et al. (2010); Ye et al. (2018); Heazlewood and Softley (2021), and tests of fundamental physics through precision measurements Andreev et al. (2018); Cairncross et al. (2017); Kozyryev and Hutzler (2017); Hudson et al. (2011); Kondov et al. (2019); Collaboration et al. (2014); Fitch et al. (2021); Cairncross et al. (2017); Yu and Hutzler (2021); Hutzler (2021); O’Rourke and Hutzler (2019); Aggarwal et al. (2018); Uzan (2003); DeMille et al. (2008); Chin et al. (2009); Kajita (2009); Beloy et al. (2010); Jansen et al. (2014); Daprà et al. (2016); Norrgard et al. (2019); Kobayashi et al. (2019); Chupp et al. (2019); Kozyryev et al. (2021). Particularly intriguing for these endeavors are polar molecules since they offer a long-range interaction due to their electric dipole moment, which leads to interaction ranges that are 3-4 orders of magnitude larger than those of the strongest magnetic atoms Pérez Ríos . Thus, it is of importance for the development of novel applications to have an accurate knowledge of the molecular properties of interest. However, this is often not the case, even for diatomic molecules. For instance, no experimental data is available on the dipolar interaction strength of aluminum monochloride (AlCl). In fact, the question of the dipole moment of AlCl was first addressed by Lide in 1965 but was considered too complex to analyze, and only a range of $1-2$ D was assigned Lide (1965). Consequently, any subsequent study used the average value of $1.5$ D as a best-guess substitute in their analyses.
In this work, we report the measurement and ab initio calculations of the permanent electric dipole moment (PDM) of AlCl. AlCl is a promising candidate for cold molecule applications that require dense molecular samples Daniel et al. (2023). Our analysis characterizes its interaction strength, which is required to accordingly develop, e.g. , experiments with AlCl in optical lattices in the future. Our comparison with ab initio theory determined that the inclusion of spin-orbit interaction is required to correctly describe AlCl. Moreover, our results differ by $\approx\!13$ from the previously used substitute value, which directly feeds into astrophysical models of stellar evolution and changes the assumed limits for the observability of AlCl in distance stars, as discussed in the following.
Both Al and Cl are created in stars through nucleosynthesis processes—aluminium predominantly through fusion reactions in high-mass stars, and Cl as a byproduct of processes such as the neon burning stage Andreazza et al. (2018); Chen et al. (2024); Jones and Nyman (2024). In the cooler outer layers of stars, most particularly the atmospheres of asymptotic giant branch (AGB) stars, these elements can combine to form AlCl in regions where temperatures and pressures allow molecular bonding Tsuji (2021); Sharp and Huebner (1990). AGB stars are in a late stage of stellar evolution, characterized by significant mass loss and complex chemistry in their outer layers Bloecker (1995); Cernicharo et al. (2015); Kane (2023). The detection of AlCl in these stars therefore provides clues toward stellar formation, chemical evolution, mass loss processes, and the distribution of elemental abundances Asplund et al. (2021). The most notable transitions to detect AlCl include the rotational transitions near 277 GHz and vibrational transitions around 12.7 microns in the infrared spectrum. Instruments like ground-based telescopes with infrared capabilities or space-based observatories have been able to detect these lines, confirming the existence of AlCl in stellar atmospheres Cernicharo and Guelin (1987); Decin et al. (2017). Furthermore, the ratio of Al to Cl can offer insights into the star’s nucleosynthesis history and the nature of the material from which the star formed. Spectroscopic monitoring of stars via precision radial velocity surveys also provide vast datasets to establish possible time variability of such spectral features Fischer et al. (2016); Halverson et al. (2016); Rubenzahl et al. (2023). However, a complete analysis of AlCl in stellar spectra has previously been hampered by limited data concerning collisional excitation states and the dipole moment Ford et al. (2004).
The relevance of AlCl extends beyond the death of stars and into the formation of new planetary systems. In circumstellar disks around young stars, Al and Cl are key components of the dust and gas that coalesce into planets Ziurys (2021). The detection of AlCl in these environments provides evidence of the chemical diversity and complexity in the material that forms planets Agúndez et al. (2012). In particular, Al is a refractory element, meaning it condenses into solid particles at high temperatures, making it a key constituent of rocky planets and asteroids Anders and Grevesse (1989). Understanding the chemical forms of Al, including its molecular bonds like AlCl, can help explain how elements are distributed in the early stages of planetary formation Lee et al. (2018).
In the atmospheres of hot exoplanets, AlCl may play a role in the high-temperature chemistry, particularly in the interaction between metals and halogens. Though Al tends to condense out at lower temperatures, hot planetary atmospheres can result in the formation of AlCl that remains in a gaseous phase. This can yield information for the vertical temperature gradients in exoplanet atmospheres, such as stratification in the atmosphere, where different chemical species are confined to specific temperature zones Fortney and Nettelmann (2010); Heng and Showman (2015); Madhusudhan (2019). For example, WASP-43b and HD 189733b are prime candidates for the detection of AlCl, given their high temperatures and the presence of other metal species Bouchy et al. (2005); Hellier et al. (2011); Inglis et al. (2024); Teinturier et al. (2024). The Mid-Infrared Instrument (MIRI) on the James Webb Space Telescope (JWST) can cover the wavelength ranges necessary to detect the vibrational transitions of AlCl, which are expected to occur around 10–15 $\mu$m Argyriou et al. (2023). The strength of the absorption features, particularly at IR wavelengths, are sensitive to the dipole moment due to the effects on the vibrational modes. Thus, an increase in the measured dipole moment of AlCl will result in a larger expected intensity in the absorption peak of the stellar spectra. Depending on the brightness and opacity of the star (such as giant stars), this may produce a measurable effect in precision measurements acquired with facilities such as JWST.
Figure 1: Experimental setup to measure the dipole moment of AlCl. A CBGB source produces a beam of AlCl that is subject to a DC electric field. The excitation laser at 261 nm is frequency-stabilized using a wavelength meter, which itself is calibrated against an optical frequency comb at regular intervals. The beat node for the comparison to the comb is produced by overlapping the lasers in a fiber coupler (Thorlabs, TW1064R5A2B) and sending the output on a grating (Thorlabs, GR25-1210). MS: Mode-shaping optics, LPF: Low-pass filter, PBS: Polarizing beam splitter, SHG: Second-harmonic generation, DFC: Difference frequency comb, HWP: Half-wave plate, PD: Photodiode, PMT: Photomultiplier tube.
A schematic of our experimental setup to measure the dipole moment of AlCl is displayed in Fig. 1. We produce a cryogenic buffer gas beam (CBGB) Hutzler et al. (2012); Barry et al. (2011) of AlCl by ablating a powdered pressed target of Al:KCl inside a helium buffer gas cell with an aperture size of $5$ mm held at  $\approx\,3.4$ K. Further details on our CBGB source, ablation technique, and vacuum apparatus are found in references Daniel et al. (2021); Lewis et al. (2021). At a distance of $70$ cm downstream from the beam source, laser-induced fluorescence spectroscopy is performed on the $X^{1}\Sigma^{+}\leftarrow A^{1}\Pi$ transition at $261.5$ nm with a laser beam that is aligned perpendicular to the molecular beam. The fluorescence light is collected with a plano-convex lens with focal length $75$ mm and imaged onto an iris for spatial filtering before being focused onto a photomultiplier tube (Hamamatsu, H10722-04) with two plano-convex lenses of focal length $100$ mm. A colored glass bandpass filter (Thorlabs, FGUV5M) with a transmission of 70% at 260 nm is added to the imaging system to minimize residual background light.
The 261.5 nm laser light is produced in a second-harmonic generation cavity with a CLBO crystal that is pumped with a $523$ nm CW laser source (Vexlum, VALO SHG) and operated at a duty cycle of 5%, which covers the transit time of molecules in the detection region, to minimize UV-induced damage on the optical components Liu et al. (2024). The output of the cavity is collimated and guided into the vacuum chamber to interact with the molecular beam. For the measurements, a power of  $\approx\,1$ mW with a beam size of  $\approx\,1.5$ mm is used. The pump laser is frequency-stabilized by referencing the pump laser to a wavelength meter and feeding back onto the laser piezo with an Arduino micro-controller. The wavelength meter is regularly calibrated by briefly locking the pump laser to an optical frequency comb, which is disciplined to the global positioning system. In the calibration cycle, the lock is implemented by measuring the beat node of the pump laser and the comb, which is set to $23$ MHz, with the FFT module of a Red Pitaya microcontroller and feeding back onto the laser piezo. We use the frequency stabilization to the wavelength meter for the experimental scans due to their large dynamic range of GHz, which is necessary for the measurement presented here.
To measure the Stark effect, the spectroscopy of the molecular beam is carried out in the presence of an electric field. The field is produced by applying a voltage to two opposing aluminum plates of size $3\times 3$ cm that are held at a distance of $1$ cm with polyether ether ketone spacers. The molecular beam is imaged in the center between the plates, which are arranged such that the electric field is perpendicular to the beam axis. The laser beam polarization for this measurement is selected to be parallel to the electric field by means of a polarizing beam splitter. The electric field is calibrated by measuring the polarizability of the ${}^{2}S_{1/2}-^{2}P_{3/2}$ transition in potassium atoms, which are a byproduct of ablating the Al:KCl mixture target (see Supplemental Material).
In the presence of an external electric field $\mathbf{E}$, the molecular energy levels are shifted and split due to the interaction of the external field with the PDM, $\bm{\mu_{e}}$, described by the Stark Hamiltonian Brown and Carrington (2003)
$H_{s}$ $=$ $-\bm{\mu_{e}}\cdot\mathbf{E}\quad.$ (1)
The remaining Hamiltonian terms include the rovibrational and hyperfine structure terms which desribe the molecular energy levels of the free molecule. For the determination of the electric dipole moment, we diagonalize the Hamiltonians of the two electronic state manifolds separately and extract the transition energies (see Supplemental Materials). The corresponding transition frequencies are then compared with the measured spectral line frequencies at different electric field amplitudes by means of a least-square fit of the data and the model. In particular, we calculate the distance of each experimental frequency, $\nu_{k}^{\textrm{exp}}$, to the nearest theoretical transition frequency, $\nu^{\textrm{theo}}_{i}$, and sum over these distances to calculate the fit residuals,
$\Delta_{r}=\sum_{k}\textrm{min}|\nu_{k}^{\textrm{exp}}-\nu^{\textrm{theo}}_{i}(\mu_{X},\mu_{A})|\quad.$ (2)
The fit has two free parameters, the values of the PDM $\mu_{X}$ for the $X^{1}\Sigma^{+}$ and $\mu_{A}$ for the $A^{1}\Pi$ state, which are determined by minimizing $\Delta_{r}$. The molecular constants are taken from the literature (see Tab. 2 in the Supplemental Material). The measured line frequencies are determined by fitting Gaussian beam profiles to the spectral data and extracting their center frequency. We use this method to measure the PDMs of two vibrational and both electronic states in AlCl.
In a separate effort, we have carried out complementary ab initio calculations of AlCl to determine the PDMs. Fig. 2 plots the ab initio computed PDM for each of the three electronic states at the equilibrium internuclear distance ($R=r_{e}$) for several spin-orbit treatments (labeled #1 to 7, see Supplementary Material for their detailed definitions). Of particular interest is treatment #1 which ignores all spin-orbit interactions, treatment #2 which includes spin-orbit interactions using a minimal number of spin states, and treatment #7 which expands upon treatment #2 by including several more spin states. Fig. 2 shows that the PDMs for the $X^{1}\Sigma^{+}$ and $a^{3}\Pi$ states are highly sensitive to the spin-orbit treatment and explains the wide range of ab initio values reported in the literature Wan et al. (2016); Yousefi and Bernath (2018); Bala et al. (2024); Woon and Herbst (2009). We note that the negative PDM values are due to the chosen geometry convention that places the Cl on the positive $z$-axis and Al at the origin. The electronegativity of Cl is larger than Al so that the dipole points from Cl to Al (i.e., in the negative $z$ direction).
Figure 2: Ab initio calculation of the permanent dipole moment for the three AlCl electronic states $X^{1}\Sigma^{+}$, $A^{1}\Pi$, and $a^{3}\Pi$ computed using seven different spin-orbit treatments and two basis sets (AV5Z blue and AVQZ red). The different treatments range from no spin-orbit (#1), including spin-orbit (#2), to extended spin-orbit (#3-7).
Figure 3: Stark splitting of the $X^{1}\Sigma^{+}(v^{\prime\prime}=0)\leftarrow A^{1}\Pi(v^{\prime}=0)$ transitions as a function of the laser detuning and the electric field. The measured R(0) transition frequencies (black points) are overlayed with the theoretical predictions (blue (Al35Cl) and green (Al37Cl) areas) of the Stark splitting, Eq. (1), for the optimal values in Tab. 1. The error bars correspond to the width of the Gaussian fits to the data. The zero-offset detuning is set to the field-free R(0) transition frequency. Inset: Residuals plot of the theory-data comparison (Eq. (2)) as a function of the two dipole moments, $\mu_{X}$ and $\mu_{A}$. The yellow dot marks the center-of-gravity of the 10% contour and optimal value for the PDMs.
Fig. 3 shows the comparison of the model and the measured transition frequencies as a function of the applied electric field. To provide the full picture of the Stark splitting in AlCl, we plot the spectrum of the two main isotopologues, Al35Cl and Al37Cl, for the Q, R(0) and R(1) transitions. Since both the Stark split Q and the R(1) transitions are either too dense in frequency or provide weaker signals due to lower line strengths for an unambiguous line assignment, we chose to measure the R(0) transition of the more abundant isotope, which is more isolated in frequency space, for this work.
The R(0) manifold splits into three branches when subject to an external electric field. The Stark shift of the manifold is dominated by the value of $\mu_{X}$, which determines the size of the frequency shift of the rotational ground state, $J^{\prime\prime}=0$, in $X^{1}\Sigma^{+}$. On the other hand, the size of the splitting of the manifold is governed by the value of $\mu_{A}$, which determines the splitting of the rotational state, $J^{\prime}=1$, in $A^{1}\Pi$. These two effects are not completely independent of each other, which leads to a correlated error for the PDMs of the two electronic states.
With the polarization parallel to the electric field, the center branch shows the strongest intensity, as shown in the example in Fig. 4 for a field of $5.7$ kV/cm. The linewidths of the measured peaks (FWHM: $\approx 150$ MHz (center branch) and $\approx 300-400$ MHz (side branches)) correspond to the theoretical expectations of the state splittings (see Supplemental Material, Fig. 5). We note that these broad features are the dominating source of uncertainty in fitting the transition frequencies of the data.
Under the influence of an external field, the two hyperfine spins in AlCl lead to a large number of individual transitions that is impractical to plot. Instead, for the comparison of the theory and data, we show the model as a colormap in Fig. 3 by summing over individual Gaussians centered at the predicted frequencies and cutting the colormap off at an intensity of 50% for illustration purposes.
The inset in Fig. 3 shows the residuals of the data to theory comparison of the transition frequencies, according to Eq. (2). We take the center of gravity of the $10$% contour line of the residual as a conservative estimate for the best value for the PDM. We also carried out the same analysis for the first vibrational state $(v^{\prime\prime}=v^{\prime}=1)$, see Supplemental Material. A summary of our results of the four PDMs of AlCl and previous literature values are given in Tab. 1. The ab initio PDMs in this work (bold) are consistent with previously reported PDMs. The differences are due to the choice of the electronic basis set (the larger AV5Z in this work) and more importantly the level of spin-orbit treatment (the extended spin basis # 7 in this work). A notable trend in the ab initio PDMs for the X and $a\,^{3}\Pi$ states is the increasing magnitude with increased vibrational excitation. This trend is also experimentally observed for the X state. In contrast, for the A state the magnitudes of both the experimental and ab initio PDMs remain about the same for $v=0$ and $v=1$. This behavior is due to the different R dependence of the electronic dipole moments for these states (see Supplementary Material for details). Also of interest is the change in the PDMs with the level of spin-orbit treatment. For the X state, the $v=0$ and $v=1$ PDMs decrease significantly in magnitude when spin-orbit interactions are included using a minimal spin basis. When the spin basis is extended, the PDM magnitude increases becoming in better agreement the with experiment. Similar behavior is seen in the $a\,^{3}\Pi$ state. In contrast, for the A state the magnitude of the PDMs remain about the same as the spin-orbit treatment is extended.
Table 1: Electric dipole moment measurement and ab initio calculation of Al35Cl in units of Debye. The results from this work are listed in bold. All other listed values are previous calculations. The ab initio results are listed without (), with minimal () and with extended () spin-orbit coupling.
State (v=0) (v=1) (at $r_{e}$)
$X^{1}\Sigma^{+}$1.687(69) -1.769(100) - exp
-1.525 -1.631 -1.493 ab initio
-1.319 -1.405 -1.325 ab initio
-1.648 -1.723 -1.636 ab initio
- - -1.309 Wan et al. (2016)
-1.630 -1.702 -1.594 Yousefi and Bernath (2018)
- - -1.604 Bala et al. (2024)
- - -1.693 Woon and Herbst (2009)
- - 1–2 Lide (1965)
$A^{1}\Pi$1.032(96) -1.055(123) - exp
-0.891 -0.890 -0.888 ab initio
-0.905 -0.896 -0.916 ab initio
-0.895 -0.868 -0.906 ab initio
- - -0.918 Wan et al. (2016)
- - -1.084 Bala et al. (2024)
$a^{3}\Pi$2.014 -2.069 -1.995 ab initio
-1.802 -1.845 -1.785 ab initio
- - -1.760 Wan et al. (2016)
- - -1.669 Bala et al. (2024)
Figure 4: Stark splitting at $\mathcal{E}=5.7$ kV/cm of Al35Cl R(0) transition. A moving average of two points is applied to the data and the zero offset detuning is set to the field-free transition. The black lines are Gaussian fits to extract the three peak center freqencies and the blue circles are the data.
Our measurements conclude the outstanding question of the value of the electric dipole moment of AlCl. Our results are in excellent agreement with our ab initio calculations and with previous theoretical literature values, which are available for the vibrational ground state. We found that including the spin-orbit coupling using an extended number of spin states is crucial for an accurate ab initio description of AlCl. Both the PDM and $T_{e}$ values become in better agreement with experiment when this level of theory is used. The ab initio analysis presented in this work also explains the origin for the wide range of PDM values reported in the literature.
This result has also confirmed that AlCl is well suited for applications that require polar molecules and allows for the strategic development of such experiments with AlCl. To the best of our knowledge, the dipole moments of the bromide and the iodide, AlBr and AlI, have only been studied theoretically or in solution Gao and Wan (2017); Yang et al. (2016); Tourky and Rizk (1957). The lightest aluminum halide, AlF, which offers excellent properties for cold molecule applications, has been studied extensively and a dipole moment of $1.515$ D was measured Truppe et al. (2019).
[…]
Acknowledgements.
L. L. and B. H. acknowledge funding from the National Science Foundation under Grant No. 2145147. M. Jr. Aguirre acknowledges funding from the University of California, Riverside. This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-21-1-0263. B. K. K. acknowledges that part of this work was done under the auspices of the U.S. Department of Energy under Project No. 20240256ER of the Laboratory Directed Research and Development Program at Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (Contract No. 89233218CNA000001).